OpenStudy (anonymous):

Find the slope of the line that contains each pair of points.(3,1)and(9,2)

OpenStudy (anonymous):

[2-1\div9-3\] = 1/6=m or slope

OpenStudy (anonymous):

[2-1\div9-3\] = 1/6=m or slope

OpenStudy (anonymous):

the slope of a line containing the points (3,1) and (9,2) is \[m=\Delta y / \Delta x\]\[m=(2-1)/(9-3)=1/6\]